# Non-perturbative renormalization group for Higgs-like models in 4D

### v1

@@ Abstract
We recently defined a model of 2 coupled SU(2) Higgs doublets which revealed an interesting structure of renormalization group flows to 1-loop when the SU(2) is broken to U(1). In this article we compute the beta functions to 3 loops and show that the 1-loop structure of flows persists to higher orders. For SU(2) broken to U(1), we conjecture a beta function to all orders. The flows can be extended to large coupling using a strong-weak coupling duality $g\to 1/g$. One finds a line of fixed points which are new conformal field theories in 4 spacetime dimensions which are non-unitary due to negative norm states but still have real eigenvalues. We also find massless flows between 2 non-trivial fixed points, and a regime with a cyclic RG flow. For the flows between points on the critical line, we compute the anomalous dimensions of the perturbations in the UV and IR, and identify some special points where anomalous dimensions are rational numbers. The model is non-unitary since the hamiltonian is pseudo-hermitian, $H^{\dagger}={\cal K}H{\cal K}^{\dagger}$. The unitary operator ${\cal K}$ satisfies ${\cal K}^{2}=1$ and this allows a projection onto positive norm states with a unitary time evolution with positive probabilities.
@@ I Introduction
There is a conspicuous shortage of well-understood conformal field theories (CFT) in 4 spacetime dimensions that aren’t simply free fields, as for asymptotically free QCD Wilczek ; Gross . For marginal perturbations, such as standard generalizations of $\phi^{4}$ for the Higgs potential, the renormalization group (RG) flows are rather limited in scope in exactly 4 spacetime dimensions. For scalar fields, with $\phi^{4}$ interactions, using the epsilon expansion in $D=4-\epsilon$ dimensions, one can find CFT’s in 3 spacetime dimensions such as the Wilson-Fisher fixed points, which have irrational anomalous dimensions WilsonFisher . In fact, this is at the heart of the so-called hierarchy problem, namely that straightforward RG arguments would indicate a much higher mass for the Higgs boson since there is no known ultra-violet (UV) fixed point. [^2] Without a known UV completion of the Higgs sector, there may even be higher dimension operators that make the theory non-unitary. This article will eventually scrutinize the idea that all quantum field theories (QFT’s) begin and end at RG fixed point CFT’s, at least for non-unitary theories.
@@
By contrast, in 2D one has a large toolkit of algebraic structures to study CFT’s, in particular the Virasoro algebra and affine Lie algebra symmetry of WZW models BPZ ; WittenWZW ; KZ . For the purposes of this article, we should point out that in the classification of minimal models of CFT based on the Virasoro algebra, one finds many more non-unitary theories than unitary, all with real eigenvalues of the hamiltonian, and many of them have important physical applications, such as the Lee-Yang edge singularity. Here the non-unitarity is manifested in negative norm states, rather than non-hermiticity of the hamiltonian. We would also like to stress that constructions of these CFT’s often requires projecting onto unitary sub-Hilbert spaces of non-unitary theories FQS . For instance, in the Coulomb gas method one must project the Hilbert space of a free boson to the proper space of the minimal model CFT DotsenkoFateev , and this can be described as a BRST procedure BernardFelder . In this way one can describe both the unitary and non-unitary minimal models of CFT in 2D. This procedure is similar to the BRST treatment of gauge theories which consistently projects out ghosts. We should also mention that the unitary CFT’s can be understood as coset reductions of current-algebras GKO , so that current algebras play a fundamental role in the classification of CFT’s.
@@
It would certainly be nice to have such a 2D playground in 4D, and we hope this article opens such a door, however narrow it may turn out to be, based on the model we recently defined ALDolls . There we constructed some new kinds of marginal perturbations of scalar fields in 4D that have a rich structure of RG flows, at least to 1-loop. The original goal of that work was to define Higgs-like models with more interesting RG flows than the usual ones. A line of ultraviolet (UV) or infrared (IR) fixed points, and perhaps more interestingly, some cyclic RG flows were found, but based only on the 1-loop approximation. [^3] The possibility of cyclic RG flows was considered long ago in Wilson’s pioneering work on the RG KWilson . The nature of these marginal operators are such that they satisfy an interesting operator product expansion similar to current algebras in 2D KZ . These models can be defined for an arbitrary Lie algebra, however with applications to Higgs physics in mind, we specialized to ${\rm SU(2)}$. This algebraic structure leads to simple formulas for the 1-loop beta functions. Since the cyclic RG flows require flowing to infinite values of the couplings, it is necessary to check if the 1-loop behavior persists to higher orders. This is main subject of this paper, which can be viewed as Part II of the RG aspects of our previous work ALDolls . We will show that the 1-loop features do indeed persist non-perturbatively in a very interesting, non-trivial manner.
@@
As defined in ALDolls and reviewed below, the hamiltonian $H$ for our models is not hermitian, but pseudo-hermitian, which makes the model non-unitary. This is not an issue for applications to statistical mechanics in 4 spatial dimensions. For quantum mechanics in $3+1$ spacetime dimensions one needs additional positivity constraints related to unitarity. This doesn’t necessarily entail an unruly hornets’ nest if the hamiltonian has some additional algebraic structure. On the other hand, if one has merely the statement $H^{\dagger}\neq H$, then there is not even a nest to speak of. Let us mention that there has recently been renewed and growing interest in non-unitary minimal models in 2D and RG flows between them MussardoLG ; Tanaka ; KlebanovMM ; Negro . In BL38 hamiltonians satisfying $H^{\dagger}={\cal K}H{\cal K}^{\dagger}$, with $K$ unitary, were classified according to ${\cal C},{\cal P},{\cal T}$ discrete symmetries, and 38 universality classes of random hamiltonians were obtained, extending the 3-fold classifications of Wigner and Dyson based on time-reversal symmetry, and the 10-fold classifications of Altland-Zirnbauer which included particle-hole symmetry Dyson ; AltlandZirnbauer . These classes have seen many applications to open quantum systems. The models considered in this paper are a special case where
$H^{\dagger}={\cal K}\,H\,{\cal K}^{\dagger},~{}~{}~{}{\cal K}^{\dagger}{\cal K}=1,~{}~{}~{}~{}~{}{\cal K}^{\dagger}=K,~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\cal K}^{2}=1.$ (1)
@@
Let us summarize the main results and organization of this article. In the next section we review the definition of the models in ALDolls . In Section III we address the non-unitarity issue and propose how to project onto positive norm states which leads to a consistent quantum mechanics, with a unitary time evolution of positive probabilities. In Section IV we compute the RG beta functions to 3-loops using just the OPE structure for fully anisotropic SU(2) perturbations, which has 3 independent couplings. In Section V we specialize to SU(2) broken to U(1), which involves only 2 couplings. Based on the structure of the 3-loop results we propose beta functions to all orders. In Section VI we study the RG flows in the resulting 2-coupling parameter space to all orders in the couplings. This is made possible by a certain RG invariant $Q$ where RG flow trajectories are constant $Q$ contours, and also a strong-weak coupling duality $g\to 1/g$ of the beta functions. The resulting picture of RG flows presents itself as a kind of master flowchart for 2 couplings, since it includes new fixed point CFT’s, marginally relevant or irrelevant, asymptotic freedom, massless flows between non-trivial fixed points, and alas cyclic RG flows. [^4] This master flowchart is natural if one starts with a single coupling, which has a marginally relevant and irrelevant direction. Opening up this flow with an additional coupling that preserves this flow along the diagonals leads to this kind of flowchart. We refer to the figures below. For the cyclic flows, the RG period $\lambda$ is a simple function of the RG invariant $Q$, namely $\lambda=2\pi/\sqrt{Q}$. The existence of a regime of couplings where the RG is cyclic calls into question the commonly accepted paradigm that every QFT begins or ends at a fixed point, and we comment on this in that section. By taking one of the couplings to be imaginary, we find massless flows between different fixed points. The latter are 4D versions of flows between non-unitary minimal models in 2D which have received much attention recently MussardoLG ; Tanaka ; KlebanovMM ; Negro .
@@ II Definition of the models
where $\Phi^{\dagger}\Phi=\sum_{i=1,2}\Phi^{\dagger}_{i}\Phi_{i}$ etc. The free theory has global ${\rm SU(2)}\otimes{\rm SU(2)}$ symmetry since the fields $\Phi,\widetilde{\Phi}$ are independent, however the interactions introduced below will break this down to the diagonal ${\rm SU(2)}$. The quantization of this theory is standard. Consider the quantization of $\Phi$, since the same applies to $\widetilde{\Phi}$. Expand the field in terms of operators $a_{\pm},a^{\dagger}_{\pm}$:
@@
Our models will involve $\Phi^{2}\widetilde{\Phi}^{2}$ marginal interactions involving the ${\cal K}$-conjugate fields $\Phi^{\dagger_{\scriptscriptstyle k}}$ and $\widetilde{\Phi}^{\dagger_{\scriptscriptstyle k}}$:
$\Phi^{\dagger_{\scriptscriptstyle k}}(x)\equiv{\cal K}\,\Phi^{\dagger}{\cal K}=\int\dfrac{d^{3}{\bf{k}}}{(2\pi)^{3/2}}\dfrac{1}{\sqrt{2\omega_{\bf{k}}}}\left(-a_{-}({\bf{k}})e^{ik\cdot x}+a^{\dagger}_{+}({\bf{k}})e^{-ik\cdot x}\right).$ (9)
@@
$\langle\Phi^{\dagger_{\scriptscriptstyle k}}(x)\Phi(y)\rangle=-\langle\Phi(x)\Phi^{\dagger_{\scriptscriptstyle k}}(y)\rangle=\dfrac{1}{4\pi^{2}|x-y|^{2}}.$ (10)
@@
$A^{\dagger_{\scriptscriptstyle k}}\equiv{\cal K}\,A^{\dagger}{\cal K}.$ (11)
The operator $\Phi^{\dagger_{\scriptscriptstyle k}}\Phi$ is not its own hermitian conjugate, but is invariant under ${\dagger_{\scriptscriptstyle k}}$ conjugation:
$(\Phi^{\dagger_{\scriptscriptstyle k}}\Phi)^{\dagger_{\scriptscriptstyle k}}=\Phi^{\dagger_{\scriptscriptstyle k}}\Phi.$ (12)
@@
$J^{a}\equiv\Phi^{\dagger_{\scriptscriptstyle k}}\sigma^{a}\Phi=\sum_{i,j}\Phi^{\dagger_{\scriptscriptstyle k}}_{i}\sigma^{a}_{ij}\Phi_{j}~{}~{}~{}~{}~{}~{}(\sigma^{a})^{\dagger}=\sigma^{a},$ (13)
where $\sigma^{a}$ are the usual Pauli matrices:
@@
with $\epsilon^{abc}$ the completely anti-symmetric tensor with $\epsilon^{123}=1$. In the above equation $\kappa=1$, however we introduced $\kappa\neq 1$ for reasons that will become clear below. For ${\rm SU(2)}$ there is no sum over “$c$” in the equation (15) and this will simplify some calculations below. The extra minus sign for $\Phi^{\dagger_{\scriptscriptstyle k}}$ in (9) is what leads to the commutator in (16) and the resulting structure constant $f^{abc}$ in the OPE (15). As we will see, the beta functions can be determined from the above OPE.
@@
$J^{\dagger_{\scriptscriptstyle k}}=J,~{}~{}~{}~{}~{}\widetilde{J}^{\dagger_{\scriptscriptstyle k}}=\widetilde{J}.$ (19)
@@
$H^{\dagger}={\cal K}H{\cal K}~{}~{}~{}\Longrightarrow~{}~{}H^{\dagger_{\scriptscriptstyle k}}=H.$ (20)

### v2

@@ Abstract
We recently defined a model of 2 coupled SU(2) Higgs-like doublets which revealed an interesting structure of renormalization group flows to 1-loop when the SU(2) is broken to U(1). In this article we compute the beta functions to 3 loops and show that the 1-loop structure of flows persists to higher orders. For SU(2) broken to U(1), we conjecture a beta function to all orders. The flows can be extended to large coupling using a strong-weak coupling duality $g\to 1/g$. One finds a line of fixed points which are new conformal field theories in 4 spacetime dimensions which are non-unitary due to negative norm states but still have real eigenvalues. We also find massless flows between 2 non-trivial fixed points, and a regime with a cyclic RG flow. For the flows between points on the critical line, we compute the anomalous dimensions of the perturbations in the UV and IR, and identify some special points where anomalous dimensions are rational numbers. The model is non-unitary since the hamiltonian is pseudo-hermitian, $H^{\dagger}={\cal K}H{\cal K}^{\dagger}$. The unitary operator ${\cal K}$ satisfies ${\cal K}^{2}=1$ and this allows a projection onto positive norm states with a unitary time evolution with positive probabilities.
@@ I Introduction
There is a conspicuous shortage of well-understood conformal field theories (CFT’s) in 4 spacetime dimensions that aren’t simply free fields, as for asymptotically free QCD Wilczek ; Gross . A well known exception is ${\cal N}=4$ SUSY Yang-Mills theory. For marginal perturbations, such as standard generalizations of $\phi^{4}$ for the Higgs potential, the renormalization group (RG) flows are rather limited in scope in exactly 4 spacetime dimensions. For scalar fields, with $\phi^{4}$ interactions, using the epsilon expansion in $D=4-\epsilon$ dimensions, one can find CFT’s in 3 spacetime dimensions such as the Wilson-Fisher fixed points, which have irrational anomalous dimensions WilsonFisher . In fact, this is at the heart of the so-called hierarchy problem, namely that straightforward RG arguments would indicate a much higher mass for the Higgs boson since there is no known ultra-violet (UV) fixed point. [^2] Without a known UV completion of the Higgs sector, there may even be higher dimension operators that make the theory non-unitary. This article will eventually scrutinize the idea that all quantum field theories (QFT’s) begin and end at RG fixed point CFT’s, at least for non-unitary theories.
@@
By contrast, in 2D one has a large toolkit of algebraic structures to study CFT’s, in particular the Virasoro algebra and affine Lie algebra symmetry of WZW models BPZ ; WittenWZW ; KZ . For the purposes of this article, we should point out that in the classification of minimal models of CFT based on the Virasoro algebra, one finds many more non-unitary theories than unitary, all with real eigenvalues of the hamiltonian, and many of them have important physical applications, such as the Lee-Yang edge singularity. Here the non-unitarity is manifested in negative norm states, rather than non-hermiticity of the hamiltonian. We would also like to stress that constructions of these CFT’s often requires projecting onto unitary sub-Hilbert spaces of non-unitary theories FQS . For instance, in the Coulomb gas method one must project the Hilbert space of a free boson to the proper space of the minimal model CFT DotsenkoFateev , and this can be described as a BRST procedure BernardFelder , or a reduction based on the SU(2) quantum group. In this way one can describe both the unitary and non-unitary minimal models of CFT in 2D. This procedure is similar to the BRST treatment of gauge theories which consistently projects out ghosts. We should also mention that the unitary CFT’s can be understood as coset reductions of current-algebras GKO , so that current algebras play a fundamental role in the classification of CFT’s.
@@
It would certainly be nice to have such a 2D playground in 4D, and we hope this article opens such a door, however narrow it may turn out to be, based on the model we recently defined ALDolls . There we constructed some new kinds of marginal perturbations of scalar fields in 4D that have a rich structure of RG flows, at least to 1-loop. The original goal of that work was to define Higgs-like models with more interesting RG flows than the usual ones. A line of ultraviolet (UV) or infrared (IR) fixed points, and perhaps more interestingly, some cyclic RG flows were found, but based only on the 1-loop approximation. [^3] The possibility of cyclic RG flows was considered long ago in Wilson’s pioneering work on the RG KWilson . The nature of these marginal operators are such that they satisfy an interesting operator product expansion with an SU(2) Lie-algebraic structure, similar to current algebras in 2D KZ . These models can be defined for an arbitrary Lie algebra, however with applications to Higgs physics in mind, we specialized to ${\rm SU(2)}$. This algebraic structure leads to simple formulas for the 1-loop beta functions. Since the cyclic RG flows require flowing to infinite values of the couplings, it is necessary to check if the 1-loop behavior persists to higher orders. This is main subject of this paper, which can be viewed as Part II of the RG aspects of our previous work ALDolls . We will show that the 1-loop features do indeed persist non-perturbatively in a very interesting, non-trivial manner.
@@
As defined in ALDolls and reviewed below, the hamiltonian $H$ for our models is not hermitian, but pseudo-hermitian, which makes the model non-unitary. This is not an issue for applications to statistical mechanics in 4 spatial dimensions. For quantum mechanics in $3+1$ spacetime dimensions one needs additional positivity constraints related to unitarity. This doesn’t necessarily entail an unruly hornets’ nest if the hamiltonian has some additional algebraic structure. On the other hand, if one has merely the statement $H^{\dagger}\neq H$, then there is not even a nest to speak of. Let us mention that there has recently been renewed and growing interest in non-unitary minimal models in 2D and RG flows between them MussardoLG ; Tanaka ; KlebanovMM ; Negro . In BL38 hamiltonians satisfying $H^{\dagger}={\cal K}H{\cal K}^{\dagger}$, with ${\cal K}$ unitary, were classified according to ${\cal C},{\cal P},{\cal T}$ discrete symmetries, and 38 universality classes of random hamiltonians were obtained, extending the 3-fold classifications of Wigner and Dyson based on time-reversal symmetry, and the 10-fold classifications of Altland-Zirnbauer which included particle-hole symmetry Dyson ; AltlandZirnbauer . These classes have seen many applications to open quantum systems. The models considered in this paper are a special case where
$H^{\dagger}={\cal K}\,H\,{\cal K}^{\dagger},~{}~{}~{}{\cal K}^{\dagger}{\cal K}=1,~{}~{}~{}~{}~{}{\cal K}^{\dagger}={\cal K},~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\cal K}^{2}=1.$ (1)
@@
Let us summarize the main results and organization of this article. In the next section we review the definition of the models in ALDolls . In Section III we address the non-unitarity issue and propose how to project onto positive norm states which leads to a consistent quantum mechanics, with a unitary time evolution of positive probabilities. In Section IV we compute the RG beta functions to 3-loops using just the OPE structure for fully anisotropic SU(2) perturbations, which has 3 independent couplings. In Section V we specialize to SU(2) broken to U(1), which involves only 2 couplings. Based on the structure of the 3-loop results we propose beta functions to all orders. In Section VI we study the RG flows in the resulting 2-coupling parameter space to all orders in the couplings. This is made possible by a certain RG invariant $Q$ where RG flow trajectories are constant $Q$ contours, and also a strong-weak coupling duality $g\to 1/g$ of the beta functions. The resulting picture of RG flows presents itself as a kind of master flowchart for 2 couplings, since it includes new fixed point CFT’s, marginally relevant or irrelevant, asymptotic freedom, massless flows between non-trivial fixed points, and alas cyclic RG flows. [^4] This master flowchart is natural if one starts with a single coupling, which has a marginally relevant and irrelevant direction. Opening up this flow with an additional coupling that preserves this flow along the diagonals leads to this kind of flowchart. We refer to the figures below. For the cyclic flows, the RG period $\lambda$ is a simple function of the RG invariant $Q$, namely $\lambda=2\pi/\sqrt{Q}$. The existence of a regime of couplings where the RG is cyclic calls into question the commonly accepted paradigm that every QFT begins or ends at a fixed point, and we comment on this in that section. By taking one of the couplings to be imaginary, we also find massless flows between different fixed points along the critical line. The latter are 4D versions of flows between non-unitary minimal models in 2D which have received much attention recently MussardoLG ; Tanaka ; KlebanovMM ; Negro .
@@ II Definition of the models
where $\Phi^{\dagger}\Phi=\sum_{i=1,2}\Phi^{\dagger}_{i}\Phi_{i}$ etc. The free theory has global ${\rm SU(2)}\otimes{\rm SU(2)}$ symmetry since the fields $\Phi,\widetilde{\Phi}$ are independent, however the interactions introduced below will break this down to the diagonal ${\rm SU(2)}$. The quantization of this theory is standard. Consider the quantization of one component $\Phi=\Phi_{i}$ for $i=1~{}{\rm or}~{}2$, since the same applies to $\widetilde{\Phi}$. Expand the field in terms of operators $a_{\pm},a^{\dagger}_{\pm}$:
@@
Consider standard parity ${\cal P}$ and time-reversal symmetry ${\cal T}$. Since parity flips the sign of momentum:
${\cal P}a_{\pm}({\bf{k}}){\cal P}=a_{\pm}(-{\bf{k}}),~{}~{}~{}{\cal P}a^{\dagger}_{\pm}({\bf{k}}){\cal P}=a^{\dagger}_{\pm}(-{\bf{k}}),~{}~{}~{}~{}~{}~{}\Longrightarrow~{}~{}~{}~{}{\cal P}\Phi(t,{\bf{x}}){\cal P}=\Phi(t,-{\bf{x}}),~{}~{}~{}{\cal P}H_{0}{\cal P}=H_{0}.$ (9)
Time-reversal also flips the sign of ${\bf{k}}$ and only differs from ${\cal P}$ in that it is anti-unitary, namely ${\cal T}z{\cal T}=z^{*}$ for $z$ a complex number. This leads to
${\cal T}a_{\pm}({\bf{k}}){\cal T}=a_{\pm}(-{\bf{k}}),~{}~{}~{}{\cal T}a^{\dagger}_{\pm}({\bf{k}}){\cal T}=a^{\dagger}_{\pm}(-{\bf{k}})~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\cal T}\Phi(t,{\bf{x}}){\cal T}=\Phi(-t,{\bf{x}}),~{}~{}~{}~{}~{}{\cal T}H_{0}{\cal T}=H_{0}.$ (10)
This implies
${\cal P}{\cal T}a_{\pm}({\bf{k}}){\cal P}{\cal T}=a_{\pm}({\bf{k}}).$ (11)
Comparing with (11), we can interpret the operators $a_{\pm}^{\dagger}$ as creating states that are even or odd under ${\cal P}{\cal T}$. It’s important to point out that unlike ${\cal P},{\cal T}$, the operator ${\cal K}$ does not act locally on fields.
Our models will involve $\Phi^{2}\widetilde{\Phi}^{2}$ marginal interactions involving the ${\cal K}$-conjugate fields $\Phi^{\dagger_{\scriptscriptstyle\kappa}}$ and $\widetilde{\Phi}^{\dagger_{\scriptscriptstyle\kappa}}$:
$\Phi^{\dagger_{\scriptscriptstyle\kappa}}(x)\equiv{\cal K}\,\Phi^{\dagger}{\cal K}=\int\dfrac{d^{3}{\bf{k}}}{(2\pi)^{3/2}}\dfrac{1}{\sqrt{2\omega_{\bf{k}}}}\left(-a_{-}({\bf{k}})e^{ik\cdot x}+a^{\dagger}_{+}({\bf{k}})e^{-ik\cdot x}\right).$ (12)
@@
$\langle\Phi^{\dagger_{\scriptscriptstyle\kappa}}(x)\Phi(y)\rangle=-\langle\Phi(x)\Phi^{\dagger_{\scriptscriptstyle\kappa}}(y)\rangle=\dfrac{1}{4\pi^{2}|x-y|^{2}}.$ (13)
@@
$A^{\dagger_{\scriptscriptstyle\kappa}}\equiv{\cal K}\,A^{\dagger}{\cal K}.$ (14)
The operator $\Phi^{\dagger_{\scriptscriptstyle\kappa}}\Phi$ is not its own hermitian conjugate, but is invariant under ${\dagger_{\scriptscriptstyle\kappa}}$ conjugation:
$(\Phi^{\dagger_{\scriptscriptstyle\kappa}}\Phi)^{\dagger_{\scriptscriptstyle\kappa}}=\Phi^{\dagger_{\scriptscriptstyle\kappa}}\Phi.$ (15)
@@
$J^{a}\equiv\Phi^{\dagger_{\scriptscriptstyle\kappa}}\sigma^{a}\Phi=\sum_{i,j}\Phi^{\dagger_{\scriptscriptstyle\kappa}}_{i}\sigma^{a}_{ij}\Phi_{j}~{}~{}~{}~{}~{}~{}(\sigma^{a})^{\dagger}=\sigma^{a},$ (16)
where $\sigma^{a}$ are the usual hermitian Pauli matrices:
@@
with $\epsilon^{abc}$ the completely anti-symmetric tensor with $\epsilon^{123}=1$. In the above equation $\kappa=1$, however we introduced $\kappa\neq 1$ for reasons that will become clear below. For ${\rm SU(2)}$ there is no sum over “$c$” in the equation (18) and this will simplify some calculations below. The extra minus sign for $\Phi^{\dagger_{\scriptscriptstyle\kappa}}$ in (12) is what leads to the commutator in (19) and the resulting structure constant $f^{abc}$ in the OPE (18). As we will see, the beta functions can be determined from the above OPE.
@@
$J^{\dagger_{\scriptscriptstyle\kappa}}=J,~{}~{}~{}~{}~{}\widetilde{J}^{\dagger_{\scriptscriptstyle\kappa}}=\widetilde{J}.$ (22)
@@
$H^{\dagger}={\cal K}H{\cal K}~{}~{}~{}\Longrightarrow~{}~{}H^{\dagger_{\scriptscriptstyle\kappa}}=H.$ (23)